# EMMOE: A Comprehensive Benchmark for Embodied Mobile Manipulation in Open Environments

### v1

@@ 4 Experiments
GPT-4o(Achiam et al., 2023) 13.33 10.51 29.79 3.57 49.38
Gemini-1.5-Pro(Team et al., 2024a) 17.33 14.79 38.03 3.39 55.91
Qwen2-VL-7B(Wang et al., 2024b) 1.00 0.50 16.55 0.59 25.00
MiniCPM-V 2.6(Yao et al., 2024) 0.67 0.57 14.45 0.06 40.00
@@
4.1 Data Augmentation
SFT Augmentation
Previous work(Zhang et al., 2024c) has shown that a standardized data format would significantly enhance model training and evaluation. Therefore, we write a uniform script to convert the original EMMOE-100 data into fixed-format conversation data. During this process, all failed subtasks will be skipped as they are treated as junk data for the SFT dataset, and we initially obtained 930 SFT data in this way, which is still insufficient for LLM training. To expand the dataset, we use GPT-4o (Achiam et al., 2023) to regenerate text descriptions of tasks and the analysis of each subtask for three times. This approach not only enhances the diversity of instructions, allowing the LLM to adapt to different user input styles, but also helps to avoid introducing additional inaccuracy or inconsistency. Finally, we obtain 3,720 SFT data in total. The relevant code and data samples are available in Appendix F.1.
DPO Augmentation
DPO (Rafailov et al., 2024) training has a strict requirement for data format, which must include prompt, chosen and rejected. For the $i$-th subtask and its input instruction $I_{i}$, if the execution of output $O_{i}$ fails but the next output $O_{i+1}$ succeeds after re-plan, we will choose $I_{i}$ as the prompt, $O_{i}$ as the rejected and $O_{i+1}$ as the chosen. Although this approach aligns well with the concept of preference data, the proportion of re-planned data is relatively low. Thus, we utilize following methods to construct new DPO data. Order Change: We shuffle the order of successful subtasks, treating successful output $O_{i}$ as chosen and $O_{i+1}$ as rejected. This approach aims to help LLMs learn the logical relationships between subtasks, particularly the optimal sequence of actions. Action Change: To standardize the planner model’s output and reduce responses outside the action list, we replace action in subtasks with non-standard names or actions outside the available list. Model Change: To enable the LLM owns the ability to select the appropriate low-level model for a given scenario, we replace the model choice with models of the same type in the model list. As a result, we get 10,104 DPO data in total. More visualized processing flows and data samples are provided in Appendix F.2.
4.2 Model Training
We select 90 tasks from EMMOE-100 as our training tasks. Using the methods described in Section 4.1, we obtain 3,316 SFT training data and 8,984 DPO training data in total. Then we select Video-LLaVA-7B (Lin et al., 2023) as our base model and conduct a two-stage training process. In the first stage, we fine-tune the base model with a learning rate of 5e-4 on 4$\times$×NVIDIA A40. In the second stage, we align the fine-tuned model with DPO and train with a learning rate of 5e-6. To prevent catastrophic forgetting and maintain the intrinsic model capability, LoRA (Hu et al., 2021) is applied in both stages, with LoRA rank set to 128 and $\alpha$ to 256 in stage one, and LoRA rank set to 8 and $\alpha$ to 8 in stage two. More training details are listed in Appendix G.
4.3 Setup
Metrics
In addition to SR, TP, SER and SRR introduced in Section 2.3, we also choose Path Length Weighted SR (PLWSR)(Shridhar et al., 2020) as one of our evaluation metrics. PLWSR is defined as SR×$\times$(length of successful trajectory) / $max$(length of expert trajectory, length of successful trajectory) and measures the ability gap between the agent and the expert in successful trajectories.
Baselines
High Level Planner: Modular framework and communication mechanism greatly facilitate the deployment of various LMMs in our HomieBot. We select four representative LMMs as baseline planners: GPT-4o (Achiam et al., 2023), Gemini-1.5-Pro (Team et al., 2024a), Qwen2-VL-7B (Wang et al., 2024b) and MiniCPM-V 2.6 (Yao et al., 2024). GPT-4o and Gemini-1.5-Pro can be easily integrated into HomieBot after minor adjustments to format requirements. By leveraging the in-context learning abilities and providing output examples for each inference, the other two models can also be deployed in our system. Low Level Executor: We extract individual skills from M3 (Gu et al., 2022) and modify their implementations. Original skills require the initial and final states of the object. We map the object name to obtain specific background information and select the nearest object. Additionally, robotic arms will be reset after each execution to enhance the success rate. We also pass all environmental state information between executions to ensure environmental consistency. We provide more details of deployment in Appendix H.1.
Evaluation Benchmarks
@@ 4.3 Setup
4.4 Results
We begin with a general evaluation since all data are unseen to baseline models. As shown in Table 2, the DPO version of HomieBot achieves the best performance in SR, PLWSR and TP metrics and significantly surpasses the baseline models. Additionally, it is evident that for open-source models of similar size, even state-of-the-art LMMs like Qwen2-VL-7B (Wang et al., 2024b) and MiniCPM-V 2.6 (Yao et al., 2024) struggle to complete EMMOE tasks without additional training. For SER, though the DPO version still performs best, the improvement is not so obvious as in other metrics, Gemini-1.5-Pro even surpasses the SFT version. This phenomenon should be attributed to the nature of SER, which reflects the model’s ability to correctly determine when a task is completed and should be terminated. It is less influenced by format requirements and low-level executions, but relies more on the model’s inherent reasoning ability. The strong reasoning capabilities of GPT-4o (Achiam et al., 2023) and Gemini-1.5-Pro (Team et al., 2024a) enable them to effectively decide when to end a trajectory.
However, SFT version performs best rather than DPO version for SRR. Since SRR reflects the model’s ability to adapt to environments and adjust from failure, we think this could be relevant with the limitations of the DPO method (Xu et al., 2024). Although DPO brings unparalleled advantages in training efficiency and convenience, it compromises the model’s generalization and transferability to certain extent. Therefore, we further evaluate HomieBot separately in training and test set. As we can observe in Table 3, while DPO version performs best on all metrics in the training split, it only outperforms SFT version on SER in the test split. Additionally, DPO version shows a significant decline on SRR and even performs slightly worse than GPT-4o and Gemini-1.5-Pro. This observation further confirms that the DPO method introduces certain generalization issues. Notably, SER remains stable for both versions across the training and test splits, which further demonstrates that SER is more related to the model’s inherent judgment ability, and our specialized handling of $End$ during dataset construction has enhanced this ability (See in Appendix F.2).
@@ 4.4 Results
4.5 Analysis
To further explore the reasons for the overall low success rate and demonstrate how HomieBot can be used to simultaneously evaluate both HLP and LLE, we conduct a detailed analysis based on the results in Section 4.4. Using the error classification in Section 3.3 and the recorded feedback, we collect all errors occurred during experiments. To identify which errors are acceptable and solvable and which are the primary causes of failure, we further classify the collected errors according to whether they appear in successful or failed trajectories, the results are shown in Figure 3.
Error Analysis
Except for $E1$ and $E2$ error that come from low-level models, each error type corresponds to different capabilities of LMMs. In failed trajectories, the predominant error type across all baseline models is $F2$ error and greatly exceeds other types. This suggests that the primary obstructive factors are physical grounding failures and model hallucinations. In practical execution processes, we observe that even models are already told the object doesn’t exist or can’t be recognized, they may still produce inappropriate outputs or repeat mistakes after several steps. This issue has been significantly improved in our models, which also highlights the significance of LMM-trainable format data. With a small amount of data combined with our data augmentation methods, LMM can build up a general understanding of the current environment, enabling outputs to be grounded and compatible with low-level models.
Besides, the proportion of failed executions for two open-source models is relatively low, indicating that most subtasks are successfully completed, which seems to conflict with the very poor SR. Based on our observations, since EMMOE includes numerous complex and long-horizon tasks, execution histories often become lengthy. When the model’s understanding ability is insufficient, it may fail to fully understand or even forget previous execution contents, ultimately resulting in meaningless outputs. Although these subtasks can be successfully executed, they contribute nothing to the final task, and even worse, they will consume remaining steps and fasten task termination. In successful trajectories, the most common error is $D1$ error. This indicates that even when the model’s spatial perception ability is insufficient, it can be adjusted through feedback information. Typically, after a $D1$ error occurs, the model will output $Go\ to$ action based on the feedback, effectively resolving this error. We conduct more detailed case study in Appendix I.
@@ 4.5 Analysis
Task Performance
We also evaluate SR for each type of task defined in Section 2.2. As shown in Table 5, typical tasks are relatively easy due to straightforward processes and fewer overall steps. The most challenging are open-ended tasks, which usually have a very long total step count, with flexible processes and results, demanding powerful capabilities from both HLP and LLE models.
5 Discussions
Limitations
Limited actions and available space in Habitat restrict the scope of task design. Besides, while standardized output enables uniform evaluation for different models, it will sacrifice certain information precision. The growing number of model inferences will also lead to additional time costs. Moreover, disparities among different simulators, simulators and the real world, pose significant challenges for the transfer and generalization of low-level models, a more general and realistic evaluation platform is needed.
Table 5: Average success rate for each type of task. The highest value of each model is highlighted in bold.
Model Short- Horizon Long- Horizon Open- Ended Logical Human- Style
HomieBot (SFT) 43.75 24.60 18.52 34.01 25.24
HomieBot (SFT+DPO) 41.67 28.11 15.38 35.86 27.88
Future Works
Real-world data collection and deployment of Homie is possible since the collection methods and evaluation metrics are independent of simulators. Besides, how to efficiently utilize historical images and videos rather than text only to optimize the memory mechanism still needs to be explored. Finally, exploring how Homie can interact or collaborate with family members to complete daily tasks is also worthful and necessary.
6 Conclusion
In our work, we first introduce EMMOE, the first unified benchmark designed to evaluate both high-level planners and low-level policies. Then we present the collection and features of EMMOE-100 and propose three novel metrics to complement existing evaluation methods. Next, we introduce our HomieBot and illustrate how its two main components HLP and LLE function. In experimental parts, we demonstrate how to use original EMMOE data to construct LMM-trainable SFT and DPO datasets and evaluate different levels of models. Finally, we conduct an in-depth analysis based on the detailed error information.
Impact Statement
This research utilizes publicly accessible models and simulators, ensuring that all data comply with privacy regulations and have been anonymized when necessary. We acknowledge the potential biases in automated data generation, as well as the risks associated with data misuse, and we have implemented appropriate measures and safeguards. This paper presents work whose goal is to advance the field of Machine Learning. To support reproducibility and further research, all codes and models will be open-source.
@@ A.1 Embodied Tasks and Benchmarks
As embodied agents and LLMs develop rapidly, many embodied tasks and benchmarks have emerged. In Embodied Question Answering (EQA) tasks, EQA-v1 (Das et al., 2018), VirtualHome (Puig et al., 2018), MT-EQA (Yu et al., 2019), MP3D-EQA (Wijmans et al., 2019), IQUAD V1 (Gordon et al., 2018), OpenEQA (Majumdar et al., 2024), HM-EQA (Ren et al., 2024a), S-EQA (Dorbala et al., 2024) contains a variety of task range to evaluate logical reasoning abilities of LLMs. BLINK (Fu et al., 2024a) for visual perception abilities of LMMs. In Vision-and-Language Navigation (VLN) tasks, R2R (Anderson et al., 2018), R4R (Jain et al., 2019) and VLN-CE (Krantz et al., 2020), SOON (Zhu et al., 2021) evaluate LLM’s capabilities under different navigation settings. ALFRED (Shridhar et al., 2020) Behavior series (Srivastava et al., 2022; Li et al., 2023a) focus on interactive household tasks OVMM (Yenamandra et al., 2023) involves picking and placing any object in unseen environments. VLA-3D (Zhang et al., 2024b) for 3D semantic scene understanding and navigation. Common manipulation datasets include MT-Opt (Kalashnikov et al., 2021), VIMA (Jiang et al., 2022), ManiSkill2 (Gu et al., 2023), Calvin (Mees et al., 2022), BridgeData-v2 (Walke et al., 2023), RH20T (Fang et al., 2023a), Open-X (O’Neill et al., 2024), AgiBot World (contributors, 2024). In mobile manipulation, RT series (Brohan et al., 2022; Zitkovich et al., 2023) and Mobile ALOHA (Fu et al., 2024b) exhibit strong capabilities. GRUTOPIA (Wang et al., 2024a) takes human participation into account. Additionally, some benchmarks focus on cross-embodiments, like RoboMIND (Wu et al., 2024). Despite numerous benchmarks, a unified benchmark and relevant task is still missing. Traditional mobile manipulation uses IL to learn entire trajectories, complicating the evaluation of intermediate processes. In our work, we propose EMMOE, a holistic benchmark designed to assess both final results and the execution process.
@@ A.2 LLMs For Robotics
LLM-driven embodied agents represent cutting-edge advancements in robotics. SayCan (Ahn et al., 2022), Palm-E (Driess et al., 2023), LLM-Planner (Song et al., 2023) and EmbodiedGPT (Mu et al., 2024) combine LLMs with complex embodied tasks. TAPA (Wu et al., 2023) and SayPlan (Rana et al., 2023) use visual modules for multi-room settings. Voyager (Wang et al., 2023), Smallville (Park et al., 2023) and Octopus (Yang et al., 2023a) use LLMs to choose pre-defined functions. L3MVN (Yu et al., 2023), ESC (Zhou et al., 2023), SayNav (Rajvanshi et al., 2023) and VLFM(Yokoyama et al., 2024) build frontier or semantic maps to navigate. ViNT (Shah et al., 2023) and NoMaD (Sridhar et al., 2024) focus on image navigation, PixNav (Cai et al., 2024) uses LLM to select target image pixel. GOAT (Chang et al., 2023) is a comprehensive navigation system. Navid (Zhang et al., 2024e) and Uni-Navid (Zhang et al., 2024d) focus on end-to-end navigation models. RT-2 (Zitkovich et al., 2023) is the first Visual Language Action (VLA) model. RoboFlamingo (Li et al., 2023b) and OpenVLA (Kim et al., 2024) are open-source VLA models. Leo (Huang et al., 2024a) focuses on multiple QA problems. Manipulate Anything (Duan et al., 2024) and Octo (Team et al., 2024b) are light models for arm control. ALOHA (Zhao et al., 2023) improves action prediction through action chunking. OPF (Peng et al., 2024) explores robust Tracking with interactive robots. RoboAgent (Bharadhwaj et al., 2024) enhances object detection and generalization, and LCB (Shentu et al., 2024) uses LLMs to generate implicit strategy goals. ManipLLM (Li et al., 2024b), VoxPoser (Huang et al., 2023), Rekep (Huang et al., 2024b) combine environmental perception and task execution.

### v2

@@ 4 Experiments
4.1 Data Augmentation
SFT Augmentation
Previous work[34] has shown that a standardized data format would significantly enhance model training and evaluation. Therefore, we write a uniform script to convert the original EMMOE-100 data into fixed-format conversation data. During this process, all failed subtasks will be skipped as they are treated as junk data for the SFT dataset, and we initially obtained 930 SFT data in this way, which is still insufficient for LLM training. To expand the dataset, we use GPT-4o [35] to regenerate text descriptions of tasks and the analysis of each subtask for three times. This approach not only enhances the diversity of instructions, allowing the LLM to adapt to different user input styles, but also helps to avoid introducing additional inaccuracy or inconsistency. Finally, we obtain 3,720 SFT data in total. The relevant code and data samples are available in Appendix F.1.
DPO Augmentation
DPO [24] training has a strict requirement for data format, which must include prompt, chosen and rejected. For the $i$-th subtask and its input instruction $I_{i}$, if the execution of output $O_{i}$ fails but the next output $O_{i+1}$ succeeds after re-plan, we will choose $I_{i}$ as the prompt, $O_{i}$ as the rejected and $O_{i+1}$ as the chosen. Although this approach aligns well with the concept of preference data, the proportion of re-planned data is relatively low. Thus, we utilize following methods to construct new DPO data. Order Change: We shuffle the order of successful subtasks, treating successful output $O_{i}$ as chosen and $O_{i+1}$ as rejected. This approach aims to help LLMs learn the logical relationships between subtasks, particularly the optimal sequence of actions. Action Change: To standardize the planner model’s output and reduce responses outside the action list, we replace actions in subtasks with non-standard names or actions outside the available list. Model Change: To enable the LLM to own the ability to select the appropriate low-level model for a given scenario, we replace the model choice with models of the same type in the model list. As a result, we get 10,104 DPO data in total. More processing flows and data samples are provided in Appendix F.2.
4.2 Model Training
We select 90 tasks from EMMOE-100 as our training tasks. Using the methods described in Section 4.1, we obtain 3,316 SFT training data and 8,984 DPO training data in total. Then we select Video-LLaVA-7B [28] as our base model and conduct a two-stage training process. In the first stage, we fine-tune the base model with a learning rate of 5e-4 on 4$\times$×NVIDIA A40. In the second stage, we align the fine-tuned model with DPO and train with a learning rate of 5e-6. To prevent catastrophic forgetting and maintain the intrinsic model capability, LoRA [36] is applied in both stages, with LoRA rank set to 128 and $\alpha$ to 256 in stage one, and LoRA rank set to 8 and $\alpha$ to 8 in stage two.
4.3 Setup
Metrics
In addition to SR, TP, SER and SRR introduced in Section 2.3, we also choose Path Length Weighted SR (PLWSR)[27] as one of our evaluation metrics. PLWSR is defined as SR×$\times$(length of successful trajectory) / $max$(length of expert trajectory, length of successful trajectory) and measures the ability gap between the agent and the expert in successful trajectories.
Baselines
High Level Planner: In baseline planner selections, GPT-4o [35] and Gemini-1.5-Pro [37] are the most popular and common closed-source models. The reasoning model o1 [38] is famous for its powerful reasoning abilities and can effectively handle complex inference tasks. As for open-source models, we choose Qwen2-VL-7B [39] and MiniCPM-V 2.6 [40], which perform well in various multi-modal tasks and have similar model sizes to HomieBot. GPT-4o, Gemini-1.5-Pro and o1 can be easily integrated into HomieBot after minor adjustments to format requirements. By leveraging the in-context learning abilities and providing output examples for each inference, the other two models can also be deployed in our system. Low Level Executor: We extract individual skills from M3 [29] and modify their implementations. Original skills require the initial and final states of the object. We map the object name to obtain specific background information and select the nearest object. Additionally, robotic arms will be reset after each execution to enhance the success rate. We also pass all environmental state information between executions to ensure environmental consistency. We use single NVIDIA A40 to run both models and provide more details in Appendix H.1.
Evaluation Benchmarks
4.4 Results
We begin with a general evaluation since all data are unseen to baseline models. As shown in Table 2, the DPO version of HomieBot achieves the best performance in SR, PLWSR, TP and SER metrics. o1 also demonstrates excellent performance in our tasks and surpasses the SFT version in some metrics. Additionally, it is evident that for open-source models of similar size, even state-of-the-art LMMs like Qwen2-VL-7B [39] and MiniCPM-V 2.6 [40] struggle in EMMOE tasks without additional training.
For SER, though the DPO version still performs best, the improvement is not so obvious. This phenomenon should be attributed to the nature of SER, which reflects the model’s ability to correctly determine when a task is completed and should be terminated. It is less influenced by format requirements and low-level executions, but relies more on the model’s inherent reasoning ability. The strong reasoning capabilities of GPT-4o [35], Gemini-1.5-Pro [37] and o1 [38] enable them to effectively decide when to end a trajectory. Moreover, as we observed during experiments, o1 is more likely to “give up” compared with other models. After several failed attempts, o1 would judge that the current task is infeasible and directly terminate task execution, while other models would continue to explore. This tendency results in relatively lower SER, and also affects TP to some extent.
For SRR, o1 performs best and SFT version performs better than DPO version. Since SRR reflects the model’s ability to adapt to environments and adjust from failure, this result suggests that o1 can better leverage feedback information to make more effective re-plans. Besides, it could also be relevant with the limitations of the DPO method [41]. Although DPO brings unparalleled advantages in training efficiency and convenience, it compromises the model’s generalization and transferability to certain extent. Therefore, we further evaluate HomieBot separately in training and test set. As we can observe in Table 3, while DPO version performs best on all metrics in the training split, it only outperforms SFT version on SER in the test split. Additionally, DPO version shows a significant decline on SRR. This observation further confirms that the DPO method introduces certain generalization issues.
Notably, SER remains stable for both versions across the training and test splits, which further demonstrates that SER is more related to the model’s inherent judgment ability, and our specialized handling of $End$ during dataset construction has enhanced this ability (See in Appendix F.2).
@@ 4.4 Results
Qwen2-VL-7B[39] 1.00 0.50 16.55 0.59 25.00
MiniCPM-V 2.6[40] 0.67 0.57 14.45 0.06 40.00
GPT-4o[35] 13.33 10.51 29.79 3.57 49.38
Gemini-1.5-Pro[37] 17.33 14.79 38.03 3.39 55.91
o1[38] 28.67 24.11 44.52 13.80 38.57
@@
4.5 Analysis
To further explore the reasons for the overall low success rate and demonstrate how HomieBot can be used to simultaneously evaluate both HLP and LLE, we conduct a detailed analysis based on the results in Section 4.4. Using the error classification in Section 3.3 and the recorded feedback, we collect all errors that occurred during experiments. To identify which errors are acceptable and solvable and which are the primary causes of failure, we further classify the collected errors according to whether they appear in successful or failed trajectories, the results are shown in Figure 3.
Error Analysis
Except for $E1$ and $E2$ error that come from low-level models, each error type corresponds to different capabilities of LMMs. In failed trajectories, the predominant error type across all baseline models is $F2$ error. This suggests that the primary obstructive factors are physical grounding failures and model hallucinations. In practical execution processes, we observe that even models are already told the object doesn’t exist or can’t be recognized, they may still produce inappropriate outputs or repeat mistakes after several steps. This issue has been significantly improved in our models, which also highlights the significance of LMM-trainable format data. With a small amount of augmented data, LMM can build up a general understanding of current environments, enabling outputs to be grounded and compatible with low-level models.
Besides, the proportion of failed executions for two open-source models is relatively low, indicating that most subtasks are successfully completed, which seems to conflict with the very poor SR. Based on our observations, since EMMOE includes numerous complex and long-horizon tasks, execution histories often become lengthy. When the model’s understanding ability is insufficient, it may fail to fully understand or even forget previous execution contents, ultimately resulting in meaningless outputs. Although these subtasks can be successfully executed, they contribute nothing to the final task, and even worse, they will consume remaining steps and fasten task termination. In successful trajectories, the most common error is $D1$ error. This indicates that even when the model’s spatial perception ability is insufficient, it can be adjusted through feedback information. Typically, after a $D1$ error occurs, the model will output $Go\ to$ action based on the feedback, effectively resolving this error. We conduct more detailed case studies in Appendix I.
@@ 4.5 Analysis
5 Conclusions and Discussions
We propose EMMOE, the first unified benchmark designed to evaluate both high-level planners and low-level policies. Then we present the collection and features of EMMOE-100 and propose three novel metrics to complement existing evaluation methods. Next, we introduce our HomieBot and illustrate how its two main components HLP and LLE function. In experiments, we demonstrate how to construct LMM-trainable SFT and DPO datasets and evaluate different levels of models. Finally, we conduct an in-depth analysis based on the detailed error information.
Limitations and Future Work
Limited actions and available space in Habitat restrict the scope of task design. Besides, standardized output will sacrifice certain information precision. The growing number of model inferences will also lead to additional time costs. Moreover, we choose to conduct evaluations solely in simulation, where all researchers are required to make assessments under the same conditions, thus ensuring optimal fairness and consistency. In the future, we’ll collect more tasks, design a more efficient system, and explore real-world evaluations.
@@ A.1 Embodied Tasks and Benchmarks
As embodied agents and LLMs develop rapidly, many embodied tasks and benchmarks have emerged. In Embodied Question Answering (EQA) tasks, EQA-v1 [42], VirtualHome [43], MT-EQA [44], MP3D-EQA [45], IQUAD V1 [46], OpenEQA [47], HM-EQA [48], S-EQA [49] contains a variety of task range to evaluate logical reasoning abilities of LLMs. BLINK [50] for visual perception abilities of LMMs. In Vision-and-Language Navigation (VLN) tasks, R2R [51], R4R [52] and VLN-CE [53], SOON [54] evaluate LLM’s capabilities under different navigation settings. ALFRED [27] Behavior series [55, 56] focus on interactive household tasks OVMM [57] involves picking and placing any object in unseen environments. VLA-3D [58] for 3D semantic scene understanding and navigation. Common manipulation datasets include MT-Opt [59], VIMA [60], ManiSkill2 [61], Calvin [62], BridgeData-v2 [63], RH20T [64], Open-X [65], AgiBot World [66]. In mobile manipulation, RT series [67, 68] and Mobile ALOHA [69] exhibit strong capabilities. GRUTOPIA [70] takes human participation into account. Additionally, some benchmarks focus on cross-embodiments, like RoboMIND [71]. EmbodiedBench [72] try to evaluate LMMs together in high-levels and low-levels. Despite numerous benchmarks, a unified benchmark and relevant task is still missing. Traditional mobile manipulation uses IL to learn entire trajectories, complicating the evaluation of intermediate processes. In our work, we propose EMMOE, a holistic benchmark designed to assess both final results and the execution process.
@@ A.2 LLMs For Robotics
LLM-driven embodied agents represent cutting-edge advancements in robotics. SayCan [73], Palm-E [11], LLM-Planner [74] and EmbodiedGPT [75] combine LLMs with complex embodied tasks. TAPA [12] and SayPlan [76] use visual modules for multi-room settings. Voyager [77], STEVE  [78], Smallville [79] and Octopus [80] use LLMs to choose pre-defined functions. L3MVN [81], ESC [82], SayNav [83] and VLFM[84] build frontier or semantic maps to navigate. ViNT [85] and NoMaD [32] focus on image navigation, PixNav [33] uses LLM to select target image pixel. GOAT [86] is a comprehensive navigation system. Navid [22] and Uni-Navid [87] focus on end-to-end navigation models. RT-2 [68] is the first Visual Language Action (VLA) model. RoboFlamingo [88] and OpenVLA [89] are open-source VLA models. Leo [90] focuses on multiple QA problems. Manipulate Anything [91] and Octo [31] are light models for arm control. ALOHA [92] improves action prediction through action chunking. RoboAgent [93] enhances object detection and generalization, and LCB [94] uses LLMs to generate implicit strategy goals. ManipLLM [95], VoxPoser [96], Rekep [18] combine environmental perception and task execution.